Solve for what is being asked.



What is the molar mass of the solute in a solution made by dissolving 34.41



g of an unknown nonelectrolyte compound in 280.0 g benzene. The boiling point of



the solution is 93.2 °C. Benzene has a boiling point of 80.1 °C and molal boiling-



point elevation constant of 2.53 °C/m.

1
Expert's answer
2022-04-18T17:27:03-0400

"\Delta T=K_b*m"


where Kb is the molal boiling-point elevation constant and m is molality


"93.2 - 80.1 = 2.53 * m => m = 5.178\ mol/kg"


"280.0\ g=0.28\ kg"


"m=\dfrac{Moles \ of\ solute}{Mass\ of\ solvent} => n(solute) =m*mass(benzene)=5.178*0.28=1.45\ mol"


"M_r(solute) = \dfrac{Mass\ of\ solute}{Moles\ of\ solute}=\dfrac{34.41\ g}{1.45\ mol}=23.73\ g/mol"